Question 5 5 pts Two objects that may be considered point masses are initially separated by a distance d. The separation distance is then increased to 3d. How does the gravitational force between these two objects change as a result of the decrease? OThe force will be one third of the initial value. O The force will be one ninth of the initial value O The force will not change since it is only dependent on the masses of the objects. The force will be three times larger than the initial value. O The force will be nine times larger than the initial value.
Answer #1

Gravitational force,

Fg = GmM / r^2

F \propto 1 / r^2

When r = d,

F = GmM / d^2

rnew = 3d

so Fnew = F / (3)^2

Fnew = F / 9

Force will be one ninth of initial value.